OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1 = −12 and a27 = 66.

OpenStudy (anonymous):

\[a_{27}=-12+26\times d=66\] solve for \(d\) first

OpenStudy (anonymous):

So, I got a27=14xd=66

OpenStudy (anonymous):

no

OpenStudy (anonymous):

solve \[-12+26d=66\] first just like you solve any linear equation

OpenStudy (anonymous):

d=3

OpenStudy (anonymous):

add \(12\) to both sides, then divide both sides by \(26\) to solve for \(d\)

OpenStudy (anonymous):

yeah \(d=3\)

OpenStudy (anonymous):

Do we plug it back again?

OpenStudy (anonymous):

so the 42 term is \[-12+41\times 3\]

OpenStudy (anonymous):

111?

OpenStudy (anonymous):

idk i didn't do it hold on

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

yeah looks good